OpenStudy (anonymous):

Can someone check this for me? its functions f(x)⋅g(x)=4x⋅(x+5) =4x⋅x+4x⋅5 =4x_2+20x =8x=20x

OpenStudy (freckles):

4xx equal to 4x^2 not 8x

OpenStudy (anonymous):

f(x)⋅g(x)=4x⋅(x+5) =4x⋅x+4x⋅5 =4x_2+20x =16x=20x

OpenStudy (freckles):

again 4xx is 4x^2 not 8x or 16x

OpenStudy (freckles):

you are trying to do more work when there is no more work to do

OpenStudy (freckles):

4x^2 and 20x is far as you can write the terms 4x^2+20x <--nothing else can be done

OpenStudy (anonymous):

okay